Circular No. 6454

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996aq IN NGC 5584 S. Nakano, Sumoto, Japan, reports that Masakatsu Aoki, Tsukioka-

cho, Toyama, has discovered a supernova (mag 15.7–16.1) in NGC 5584 on CCD frames taken on Aug. 17.487 UT with a 0.43-m f/6 reflector. Aoki measured the position of the supernova: α = 14^{h}22^{m}22^{s}.73, δ = −0^{o}23′24′′.3 (equinox 2000.0), which is 15′′ west and 8′′ south of the center of NGC 5584. No star was present at this position on frames taken by Aoki on Apr. 28, May 14, 19, and 26 or on frames taken by R. Kushida (Yatsugatake South Base Observatory) on Apr. 18 and May 24. R. and Y. Kushida report the following position end figures and magnitude for SN 1996aq from an unfiltered CCD frame taken on Aug. 18.444: 22^{s}.72, 23′′.8, mag 15.0.

S. Benetti, M. Turatto, and T. Augusteijn, European Southern Observatory (ESO); L. Reduzzi, Universitá di Milano; and A. Clocchiatti, Cerro Tololo Interamerican Observatory, report that inspection of a preliminarilyreduced CCD spectrum (range 346–1047 nm, resolution 0.9 nm), obtained on Aug. 19.97 UT with the ESO 1.52-m telescope (+ Boller & Chivens spectrograph), confirms that SN 1996aq is a supernova of type Ic around maximum. The spectrum is dominated by lines (with P-Cyg profile) of Ca II (H and K + infrared triplet), Fe II, and Na I D. A relatively strong, symmetric emission, centered at ∼ 657.4 nm (rest frame), with FWZI ∼ 12 000 km/s, could be identified with Hα. The O I 777.4-nm emission is still weak in comparison with SN 1994I two days before maximum (Clocchiatti et al. 1996, ApJ 462, 462). The expansion velocity deduced from the minimum of the Ca II lines, corrected by the recession velocity of 1643 km/s (measured from the Hα of the parent galaxy), is ∼ 15 500 km/s. Towards the red end of the spectrum, a weak and broad absorption is present, which could be the minimum of the P-Cyg profile of He I (1083 nm) at an expansion velocity of ∼ 21 700 km/s. Infrared spectroscopic observations to confirm the presence of hydrogen and helium are urged. Magnitudes of SN 1996aq obtained on Aug. 18.98 with the Dutch 0.9-m telescope are: V = 14.7, U − B = +0.1, B − V = +0.6, V − R = +0.2, R − I = 0.0.

NOVA SAGITTARII 1996 Additional photometry from G. Sostero et al. (cf. IAUC 6453): Aug.

15.868 UT, V = 11.93, B − V = +1.02, V − R = +0.85, V − I = +1.64; 16.854, 11.81, +1.06, +0.79, +1.59.

1996 August 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green